OpenStudy (anonymous):

A basketball was thrown from one end of a court to the other. It had an initial upward velocity of 48 feet per second. How long did it take for the ball to hit the ground? Use the formula h = vt – 16t2.

OpenStudy (anonymous):

@Loujoelou Wanna help? :p

OpenStudy (anonymous):

Well in my virtual school I would have used the formula h= -16t^2+vt+s. The velocity would be 48 so -16t^2+48+0 (I assume that the starting height is 0) to find t we use t=-b/2a So we have t=-48/2(-16) and we get -48/-32 which is equal to 1.5 Now we put this into the equation h= -16(1.5)^2+48(1.5)+0 h=-36+72 h=36 That's the only thing I know how to solve for in these equations so hopefully I answered the right thing.

OpenStudy (anonymous):

A:3 seconds C:-3 seconds B:0.3 seconds D:-0.3 seconds

OpenStudy (anonymous):

Ugh lol thought it was asking for that. Well found the answer on this website- http://answers.yahoo.com/question/index;_ylt=AjKUMTSRV8XtbsqtyLvpTwUjzKIX;_ylv=3?qid=20110530112216AA42Ohx

OpenStudy (anonymous):

Cause Idon't know how to solve for the seconds so yeah.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

No prob. :)